OpenStudy (anonymous):

Write the equation of the line, with the given properties, in slope-intercept form. Slope = -2, through (-4,4)

OpenStudy (anonymous):

y - y1 = m(x-x1) m is the slope (x1,y1) is the point

OpenStudy (anonymous):

whis is y and which is x?? I think that is where most of my confusion starts

OpenStudy (anonymous):

I know that the -2 is my "m"

OpenStudy (anonymous):

u dont have to touch x or y :) just plug in the values u have

OpenStudy (anonymous):

so where would I place the -4 and the 4?

OpenStudy (anonymous):

x1 y1

OpenStudy (anonymous):

ok so I should now have this.... y-4=-2(x--4) I just knwo that I am doing something wrong that does not look correct to me.

OpenStudy (anonymous):

y-4=-2(x--4) its correct y-4=-2(x--4) y-4=-2(x+4) distribute y-4=-2x-8 add 4 to both sides y = -2x - 8 +4 y = -2x -4

OpenStudy (anonymous):

I SEE!!!! I get it!!! I copied this down in my notes! THANK YOU SO MUCH